OpenStudy (anonymous):

Sue has 7/8 of her bookleft to read.If she reads another 2/7 of the book, how much of the book will she have left? A 33/56 B 1/5 C 5/8 D 23/40

OpenStudy (anonymous):

(7/8)*(7/7)=(49/56) (2/7)*(8/8)= (16/56) So (49/56)-(16/56)= 33/56